# VideoUFO: A Million-Scale User-Focused Dataset for Text-to-Video Generation

### v1

@@ 11 Ablation Studies on Curation of VideoUFO
Figure 10: The relationship between pre-set number of clusters for K-means and final number of user-focused topics.
Table 3: The methods for clip verification. We use $1,000$ clips for this ablation study. “Direct GPT-4o” refers to processing videos directly with GPT-4o, and “Overlap” measures the extent to which the predictions of other methods match those of “Direct GPT-4o”.
  Method ##\## Verified Overlap API Cost
Direct GPT-4o $507$- \$$\$162.23$
Brief GPT-4o $174$ $0.621$$\$0.04$
GPT-4o-mini $276$ $0.657$$\$0.01$
GPT-4o $299$ $0.770$$\$0.46$
Detail GPT-4o-mini $393$ $0.806$$\$0.03$
12 The List of Full Topics
  dance horror alien cat
music fashion walk forest
cyberpunk portrait flight battle
drive nature rain sunset
dog celebration romance beach
mythology art explosion conflict
eat zombie run landscape
moon robot space village
dragon cityscape lion robotics
rabbit race desert night
christmas bird cook culture
paint girl visualization beauty
animal play fire magic
fish war jungle game
car architecture boy astronomy
coffee mountain weather warrior
snow logo wave monkey
swim friendship haunt cloud
ufo smoke garden gothic
woman cartoon chase
cinematography
flower conversation ghost angel
elephant psychedelia superheroes future
  vehicle scenery dinosaur humor
celebrity underwater sunrise history
combat ocean ritual apocalypse
emotion meditation demon bear
wildlife transportation cyborg warfare
mystery light fairy monster
motorcycle royalty tree tiger
wolf sleep family mermaid
batman futurism comedy tourism
travel panda lake water
egypt train drone interaction
astronaut horse child soccer
destruction design samurai
chinese culture
frog smile religion sing
advertise mar sky winter
butterfly fox cosmos wind
hair box aviation growth
pirate festival steampunk guitar
vampire censorship medicine mushroom
halloween couple love spaceship
shark soldier exercise eye
read river fruit martial arts
luxury snake kitten urbanization
mouse elon musk santa claus castle
spacecraft atmosphere basketball temple
indian culture
loneliness man storm
squirrel nightlife ant office
action spirituality noir star wars
iron man timelapse meme metropolis
hinduism imagery chicken survival
artificial intelligence
fall decoration performance
whisker football pig santa
princess city sadness paranormal
crime insect automobile arte
duck flag nostalgia accident
neon study waterfall tema
shop turtle viking supernatural
wed jesus concert character
cow witch interior journey
ninja dreamscape owl rap
goddess literature news boat
jump clown nighttime psychedelics
disney tranquility cannabis medieval
volcano firearm pikachu dj
storytelling filmmaking spiderman ship
monk expression
swordsmanship
rainforest
barbie skateboard piano market
marine life military model surf
ice cream school television infant
firework neuroscience creature vlogging
sculpture the matrix experiment time travel
party gold bible park
camera movement
meteor octopus whale
cycle space battle eagle autumn
din spider-man crash finance
door skyscraper launch tea
escape face
food combination
fight
arab culture cemetery room ai
  time police spider bloom
crow conversion wealth doll
nightclub phone botany apple
night sky sport trump balloon
pizza earth
cryptocurrency
crystal
parody banana rotation environment
black hole heart dolphin planet
death
haunted mansion
superman knight
dragonfly giant russia krishna
cake bunny countryside statue
politics cowboy origami gather
childhood penguin drift act
new york reflection perseverance star
dystopia cricket
horseback riding
jellyfish
routine classroom sheep rise
money climb witchcraft deity
chaos laughter fractal evolution
snowfall clothe treasure flying car
cosmology skeleton king pyramid
street glass fear candy
dream ancient rome color agriculture
prayer book meet disaster
island
virtual reality
sea creature tokyo
parent teddy bear investigation food
traffic bicycle farmer jewelry
gangster violence commercial encounter
advertisement bollywood youth teenager
symbolism pixar aesthetic solar
deer wizardry victory neighborhood
mutation crustacean library puppy
religious imagery
portal cafe crossover
china chess elf parrot
construction gameplay hell pollution
ancient greece
greet bounce ambiance
gallop sound toy ballet
robot combat attire videogame beer
fitness costume heaven islam
iceberg paris werewolf hospital
footage drum satan cat behavior
sun galaxy blink rock
mysticism simulation civilization mask
bar trade archaeology divinity
relationship farm graffiti dentistry
lionel messi employment protest assassination
toilet
slavic mythology
youtube wine
education homelessness shadow resilience
mansion bakery vegetable gamble
sail ceremony unicorn speed
marvel gorilla africa visual effects
skyline humanoid godzilla clock
clean write chocolate realism
sci-fi egg tropical thriller
yoga cuisine dark bodybuilding
mirror hack determination thunderstorm
rave turkey bridge elderly
abandonment puppetry goat appearance
minecraft render rainbow be
sitcom reunion bedroom house
  hulk anger taylor swift hybridization
hunt
communication
happiness meadow
background field skull dinner
lightning tennis candle intelligence
swimming pool
energy disco relaxation
gift business bus illness
seaside program alcohol disapproval
imprisonment kitchen gesture cry
punk hand nutrition social media
record floral life soviet
truck aerial draw shiva
compassion digital ski bubble
italy retro computer cinema
wrestle fog tribalism hallway
crocodile restaurant lego cybersecurity
giraffe
cristiano ronaldo
speech camp
sand introspection
character design
peacock
textile raccoon joy cuteness
victoriana spaghetti church manufacture
blonde living room skydive fatherhood
heroism moonlight satire helicopter
picnic bikini safety extraterrestrial
leaf motivation telephone hotel
bee spongebob occult nuclear
marine astrology teach photoshoot
tornado hug display asteroid
cherry blossoms
prehistory health robbery
aurora
music production
hip-hop darkness
tattoo brazil breakdancing darth vader
goku flood suspense geometry
digital art science student circus
emergency ganesha joker cheese
behavior stargaze theater quantum
waste physic delivery folklore
pokemon shoe holography card
mexico coral reefs machinery villain
sneaker tesla dive prophet
woodworking palace
anthropomorphism
surveillance
coin burger blue muppets
korea diplomacy calendar antique
einstein campfire presentation flame
superpower bottle grandmother bedtime
message collaboration interview drug
pixel art jazz body electricity
anatomy rescue terrorism sea monster
primate donkey brand skate
mentorship beverage dog eating lizard
burn cabin wonderland event
makeup milk archery connectivity
rebellion splash synthesizer pumpkin
eiffel tower surprise kiss contemplation
urban buddhism mathematics bully
map mecha
nuclear warfare
chemistry
wisdom tunnel lo-fi buddha
science fiction
procrastination
print farewell
genetics
mercedes-benz
indonesia fable
foot diamond secrecy duel
dynasty cell philosophy economics
  bmw metalworking strawberry biblical
golf hope abstraction weightlift
camel gratitude colonialism
bioluminescence
corgi crowd breakfast violin
cliff bat exhibition signage
dubai
fluid dynamics
morocco bag
diversity selfie empowerment worm
climate afterlife
amusement park
malfunction
geology microbiology
head movement
imran khan
cybernetics
public speaking
stock market purple
shrek intimacy sibling maya
android pond firefighting pregnancy
tsunami leadership iran mcdonald’s
confusion aztec lightsaber hydration
mosque baseball kingdom sphere
soul surgery perspective muslim
maze
mickey mouse
mario
mental health
gender espionage conspiracy muscle
techno security levitation creativity
seascape vehicle alphabet grasshopper
repair parade style tarot
shepherd migration cooky vaporwave
injury season furniture metal
potato firefly wake illusion
silhouette treehouse bath trailer
companionship
skincare terminator super mario
pakistan kungfu argument psychology
bullfighting perfume melt data
oil watermelon heal antarctica
empire earthquake university phoenix
grief himalaya entertainment wall
real estate putin discovery redhead
jugando afrofuturism soda hanuman
slavery cottage kaiju tyrannosaurus
sunshine businessman hong kong glitch
laptop diwali
street photography
hedgehog
carnival orange napoleon wash
community awaken
mortal kombat
skin
law gladiator blood justice
noah’s ark anxiety hamster spain
heist tolkien stadium bigfoot
tank charity fungus mammoth
eminem pope craft cinderella
chair noodle hygiene biomechanics
deception pandemic courtroom pug
fast food sunlight titanic hamburger
campus supercar children’s art rooster
hotdog thailand temperature
naval warfare
expressionism typography luck polar bear
charge type snail swamp
laser ethereal subterranean doraemon
riot confrontation mahabharata predation
school bus zoo ceo mongol
monroe sushi collision ink
lgbtq+ superhero goblin spartan
donut atlantis
web development
minimalism
panther iphone cryptids e-commerce
submarine lip-sync beard vend
  lava stress airport dessert
yacht particle capybara lotus
celtic prince caterpillar nurse
nebula indoors aquarium mujer
lemon
ottoman empire
fireplace warhammer
radio virus cube currency
cup examination infidelity courage
honey transformer unity connection
baldness arrest influencer proposal
lifestyle knife gas station taxi
content
metamorphosis
entrepreneurship
gandhi
sandwich slime suit raven
biology emoji shower spacewomen
dog smiling crucifixion priest nativity
minion emergence taco language
graduation ballistics inspiration alley
cleopatra kidnap hill preach
thor consultation prophecy fan
cub plague dracula motherhood
childbirth kangaroo sonic peru
businesswoman
saturn burnout juice
watch spice dunk aspiration
square umbrella automation puzzle
anticipation hallucination navigation mandala
rama search breathe wish
cannibalism plumb fairytale mine
sloth engine seagull australia
pineapple logistics blindness backroom
impact pottery mindfulness customer
alligator
decision-making
greed hockey
indoor long rollercoaster transition
snack deadpool game show volleyball
laundry cheetah alexander cocktail
existentialism sikhism imagination alchemy
scorpion squid
schwarzenegger
geography
decay massage freedom fatigue
cosplay adam driver tupac hippy
llama interracial chameleon dust
knit gundam
nursery rhyme
curse
nail patriotism handshake sniper
mughals notification profession workspace
augmented reality
pool party morality popcorn
luffy commute
shawn mendes
independence
snoop dogg contrast observation mutant
swan monastery digestion group
abraham interrogation vote seahorse
drought taj mahal comet therapy
niño teletubbies caminar jupiter
positivity ladybug sofa olympics
drink salad homecoming vase
confidence cuba remote work circulation
soup cancer stormtrooper venom
blade runner consciousness addiction gingerbread
sew
astral projection
sri lanka philippine
fry lecture peach forage
retrofuturism pastry harley quinn gravity
blacksmithing lily inquisitor rugby
landmark ramayana
genshin impact
backflip
  zelda urban decay tea ceremony toast
chaplin asmr elvis hypnosis
curtain throw marathon odyssey
persona teleportation theory temptation
kurt cobain rapunzel hatsune miku destiny
division bass space shuttle yin-yang
billie eilish
sleep paralysis
horizon garfield
biblical-figures
mia khalifa dua lipa
unreal engine
half-life dieselpunk chuck norris eggplant
tower
deepika padukone
puppeteering limitation
seinfeld oracle garvey

### v2

@@ Appendix C Ablation Studies on Curation of VideoUFO
Appendix D The Distributions of Caption Lengths and Clip Durations
Fig. 10 presents the statistical distributions of caption lengths (for both brief and detailed captions) and clip durations.
Appendix E The Distributions of Video Quality Score
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We follow this common practice for writing a paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See the Limitation at the end of Sections 4 and 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This is not a theory paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All the information needed to reproduce the main experimental results are provided in the main paper as well as the Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The dataset is publicly available at https://huggingface.co/datasets/WenhaoWang/VideoUFO, and the code is publicly available at https://github.com/WangWenhao0716/BenchUFO.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental settings and details are provided in the main paper as well as the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: All the experimental results are significant and stable, and error bars are not reported because it would be too computationally expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We include the experiments compute resources in each related section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The data used in this study is publicly available on YouTube, and we do not anticipate any negative/positive societal impacts resulting from the release of our dataset.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We only scrape data with explicit CC BY 4.0 License to avoid any potential copyright infringement.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all the assets used in our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Please see the README file in the released dataset.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve that.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve that.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We only use LLM for writing, editing, or formatting purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.